# BPP-Search: Enhancing Tree of Thought Reasoning for Mathematical Modeling Problem Solving

### v2

@@ Abstract
LLMs exhibit advanced reasoning capabilities, offering the potential to transform natural language questions into mathematical models. However, existing open-source datasets in operations research domain lack detailed annotations of the modeling process, such as variable definitions, focusing solely on objective values, which hinders reinforcement learning applications. To address this, we release the StructuredOR dataset, annotated with comprehensive labels that capture the complete mathematical modeling process. We further propose BPP-Search, a algorithm that integrates reinforcement learning into a tree-of-thought structure using Beam search, a Process reward model, and a pairwise Preference algorithm. This approach enables efficient exploration of tree structures, avoiding exhaustive search while improving accuracy. Extensive experiments on StructuredOR, NL4OPT, and MAMO-ComplexLP datasets show that BPP-Search significantly outperforms state-of-the-art methods. In tree-based reasoning, BPP-Search excels in accuracy and efficiency, enabling faster retrieval of correct solutions.
@@
Teng Wang1, Wing-Yin Yu*2, Zhenqi He1, Zehua Liu1, Xiongwei Han3, Hailei Gong3, Han Wu2 Wei Shi2, Ruifeng She2, Fangzhou Zhu3, Tao Zhong3 1Department of Mathematics, the University of Hong Kong, Hong Kong SAR, China 2Noah’s Ark Lab, Huawei, Hong Kong SAR, China, 3Noah’s Ark Lab, Huawei, Shenzhen, China, Correspondence: rocket.YuWingYin@huawei.com
@@ 2.2 Process Reward Model
Several works Cobbe et al. (2021); Lightman et al. (2023); Uesato et al. (2022) introduced the concept of the Process Reward Model (PRM), demonstrating its ability to significantly enhance the performance of weak and small-scale policy models. Compared to the Outcome Reward Model (ORM), PRM achieves better performance but incurs higher labeling costs Uesato et al. (2022). Initially, PRM training relied on manually labeled data Lightman et al. (2023). To address the growing demand for processing labels, Monte Carlo Tree Search (MCTS)-based methods Wang et al. (2024d, a); Luo et al. (2024); Zhang et al. (2024); Setlur et al. (2024) were later developed to simulate and assign scores to reasoning processes. While effective, MCTS-based approaches require wide and deep trees to generate labeled data through extensive rollouts for score convergence at intermediate nodes, resulting in extremely high computational resource demands. In contrast, manually labeled data is deterministic and directly reflects the intended reasoning process without relying on approximations.
@@ 3.2 StructuredOR Dataset Framework
Building on Xiao’s work Xiao et al. (2024), which introduces a framework for generating abstract mathematical models, we refine and expand this approach to cover a wider spectrum of abstract models. We leverage LLM such as GPT-4o Achiam et al. (2023) in conjunction with mathematical solvers to instantiate abstract models into concrete examples. Furthermore, we implement a series of validation mechanisms to construct and verify the accuracy of these concrete problems, thereby ensuring the dataset’s quality and reliability. In summary, the StructuredOR dataset provides pairs of concrete questions and corresponding model data, accompanied by comprehensive annotations detailing the entire modeling process. Fig. 3 illustrates the construction pipeline, with each step distinguished by a color-coded arrow: blue for Step 1, orange for Step 2, red for Step 3, and green for Step 4. The whole process is delineated as follows:
@@
Finally, we introduce an iterative strategy for data augmentation by changing parameter data distributions during the initial generation phase. This yields a dataset comprising 124 concrete questions and their corresponding models, spanning domains including logistics, scheduling, and networks, of which 77 examples are from the original framework, and 47 are generated through data augmentation.
@@ 3.3 PRM Dataset Preparation
MCTS-based methods rely on wide and deep trees to generate labeled process data through extensive rollouts, ensuring score convergence for intermediate nodes but demanding significant computational resources. In contrast, manually labeled data is deterministic and directly captures the intended reasoning process without relying on approximations.
@@
GPT-4o 19 19 21 21 23 30
GPT-4o-mini 13 14 19 12 19 21
Llama-3-70B 10 17 19 17 17 22
Llama-3.1-70B 18 10 21 5 14 17
Llama-3.2-11B 2 0 1 0 0 8
Qwen-2-72B-Instruct 2 2 4 3 1 5
Qwen-2.5-MATH-72B-Instruct 2 2 0 0 0 4
Qwen-2.5-72B-Instruct 6 5 9 2 2 2
Mixtral-8$\times$7B-v0.1 0 0 0 0 1 1
@@ 4.2 BPP-Search
Fig. 4 illustrates the BPP-Search process, where nodes are pruned based on PRM scores during beam search (Width = 2), and the Pairwise Preference Algorithm ranks the final candidates to select the optimal solution, ensuring robustness and accuracy.
@@ 5 EXPERIMENT
Dataset Number of Problems Resolved
StructuredOR 30/38
NL4OPT 143/289
ComplexLP 72/211
@@ 5.1 Baseline
Metric Accuracy Precision Recall F1-Score
Value 0.9823 0.9772 0.9868 0.9820
Metric Accuracy Precision Recall F1-Score
Value 0.7560 0.7761 0.7196 0.7468
Table 5: Accuracy and reasoning steps for our methods and baselines with a fixed policy model (GPT-4o) on the StructuredOR, Mamo-ComplexLP Huang et al. (2024), and NL4OPT Ramamonjison et al. (2022) test datasets. The results are based on 30 problems from StructuredOR, 72 from Mamo-ComplexLP, and 143 from NL4OPT that are confirmed solvable by the policy model in prior experiments.
@@ 5.2 Evaluation of BPP-Search
We evaluate our methods on the solvable problems identified in the datasets, as described in Section 5.1. Table 5 presents a comparison between our methods and baseline approaches, focusing on correct rate and reasoning steps. The results show that, under the condition where none of the methods fine-tune the policy model, our methods achieve superior performance with fewer reasoning steps, significantly outperforming baselines.
@@ A.1 Modeling Data Format Specification
Tables  789101112131415 define the standardized data format for representing mathematical models. This convention provides a structured and consistent way to label and organize modeling data, ensuring clarity and usability across different tasks and datasets.

### v3

@@ Abstract
LLMs exhibit advanced reasoning capabilities, offering the potential to transform natural language questions into mathematical models. However, existing open-source datasets in operations research domain lack detailed annotations of the modeling process, such as variable definitions, focusing solely on objective values, which hinders reinforcement learning applications. To address this, we release the StructuredOR dataset, annotated with comprehensive labels that capture the complete mathematical modeling process. We further propose BPP-Search, an algorithm that integrates reinforcement learning into a tree-of-thought structure using Beam search, a Process reward model, and a pairwise Preference algorithm. This approach enables efficient exploration of tree structures, avoiding exhaustive search while improving accuracy. Extensive experiments on StructuredOR, NL4OPT, and MAMO-ComplexLP datasets show that BPP-Search significantly outperforms state-of-the-art methods. In tree-based reasoning, BPP-Search excels in accuracy and efficiency, enabling faster retrieval of correct solutions. The StructuredOR dataset is available at https://github.com/tengwang0318/StructuredOR.
@@
Teng Wang1, Wing-Yin Yu2, Zhenqi He1, Zehua Liu1, Hailei Gong2, Han Wu2, Xiongwei Han2, Wei Shi2, Ruifeng She2, Fangzhou Zhu2, Tao Zhong2 1the University of Hong Kong 2Noah’s Ark Lab, Huawei, wt0318@connect.hku.hk
@@ 2.2 Process Reward Model
Several works Cobbe et al. (2021); Lightman et al. (2023); Uesato et al. (2022) introduce the concept of the Process Reward Model (PRM), demonstrating its ability to significantly enhance the performance of weak and small-scale policy models. Compared to the Outcome Reward Model (ORM), PRM achieves better performance but incurs higher labeling costs Uesato et al. (2022). Initially, PRM training relied on manually labeled data Lightman et al. (2023). To address the growing demand for processing labels, Monte Carlo Tree Search (MCTS)-based methods Wang et al. (2024d, a); Luo et al. (2024); Zhang et al. (2024); Setlur et al. (2024) were later developed to simulate and assign scores to reasoning processes. While effective, MCTS-based approaches require wide and deep trees to generate labeled data through extensive rollouts for score convergence at intermediate nodes, resulting in extremely high computational resource demands. In contrast, manually labeled data is deterministic and directly reflects the intended reasoning process without relying on approximations.
@@ 3.2 StructuredOR Dataset Framework
Set Parameter Variable Constraint
Average 2.22 3.95 1.40 2.38
Max 4 7 4 4
Standard Deviation 0.66 1.07 0.70 0.85
Table 1: Descriptive statistics for the abstract modeling components in the StructuredOR dataset. The table reports the number of sets, parameters, variables, and constraints in the abstract model (note that these numbers will expand after the model is instantiated into a concrete form).
Building on  Xiao et al. (2024)’s work, which introduces a framework for generating abstract mathematical models, we refine and expand this approach to cover a wider spectrum of abstract models. We leverage LLM such as GPT-4o Achiam et al. (2023) in conjunction with mathematical solvers to instantiate abstract models into concrete examples. Furthermore, we implement a series of validation mechanisms to construct and verify the accuracy of these concrete problems, thereby ensuring the dataset’s quality and reliability. In summary, the StructuredOR dataset provides pairs of concrete questions and corresponding model data, accompanied by comprehensive annotations detailing the entire modeling process. Fig. 3 illustrates the construction pipeline, with each step distinguished by a color-coded arrow: blue for Step 1, orange for Step 2, red for Step 3, and green for Step 4. The whole process is delineated as follows:
@@
Concrete Variables per Abstract Variable Frequency
1 70
2 214
3 36
4 143
6 5
8 22
Table 2: Distribution of the number of instantiated variables derived from each abstract variable in the concrete model.
@@
Category Agriculture Logistics Education Sports Military Energy Telecommunications Manufacturing Health Services Finance
Count 24 20 15 12 12 12 11 9 5 4
Table 3: Instance category distribution in the StructuredOR dataset, demonstrating broad applicability across industries.
Finally, we introduce an iterative strategy for data augmentation by changing parameter data distributions during the initial generation phase. This yields a dataset comprising 124 concrete questions and their corresponding models, spanning domains including logistics, scheduling, and networks, of which 77 examples are from the original framework, and 47 are generated through data augmentation. Table 1 presents the statistical distribution of the key components—sets, parameters, variables, and constraints—in the abstract model of the StructuredOR dataset. Table 2 summarizes the number of concrete variables instantiated from each abstract variable, while table 3 shows the industry distribution. Together, these tables illustrate the dataset’s structural diversity and comprehensive domain coverage.
@@ 3.3 PRM Dataset Preparation
MCTS-based methods rely on wide and deep trees to generate labeled process data through extensive rollouts, ensuring score convergence for intermediate nodes but demanding significant computational resources and incurring Reward Hacking Weng (2024). In contrast, manually labeled data is deterministic and directly captures the intended reasoning process without relying on approximations.
@@ 4.2 BPP-Search
GPT-4o 19 19 21 21 23 30
GPT-4o-mini 13 14 19 12 19 21
Llama-3-70B 10 17 19 17 17 22
Llama-3.1-70B 18 10 21 5 14 17
Llama-3.2-11B 2 0 1 0 0 8
Qwen-2-72B-Instruct 2 2 4 3 1 5
Qwen-2.5-MATH-72B-Instruct 2 2 0 0 0 4
Qwen-2.5-72B-Instruct 6 5 9 2 2 2
Mixtral-8$\times$7B-v0.1 0 0 0 0 1 1
Dataset Number of Problems Resolved
StructuredOR 30/38
NL4OPT 143/289
ComplexLP 72/211
@@
Fig. 4 illustrates the BPP-Search process, where nodes are pruned based on PRM scores during beam search (Width = 2), and the Pairwise Preference Algorithm ranks the final candidates to select the optimal solution, ensuring robustness and accuracy. Appendix A.6 analyses computational cost of BPP-Search.
@@ 5.1 Baseline
Table 6: Accuracy and reasoning steps for BPP-Search and baselines with a fixed policy model (GPT-4o) on the StructuredOR, Mamo-ComplexLP Huang et al. (2024), and NL4OPT Ramamonjison et al. (2022) test datasets. The results are based on 30 problems from StructuredOR, 72 from Mamo-ComplexLP, and 143 from NL4OPT that are confirmed solvable by the policy model in prior experiments.
@@ 5.2 Evaluation of BPP-Search
We evaluate our methods on the solvable problems identified in the datasets, as described in Section 5.1. We perform supervised fine-tuning of the PRM and Preference Model on their corresponding training datasets, as described in Sec.3.3 and Sec.4.2, respectively. Detailed evaluation results are provided in Appendix A.7. Table 6 presents a comparison between our methods and baseline approaches, focusing on correct rate and reasoning steps. The results show that, under the condition where none of the methods fine-tune the policy model, our methods achieve superior performance with fewer reasoning steps, significantly outperforming baselines.
@@ 7.1 Trade-offs in ToT Structure: Performance and Computational Cost
7.2 Dilemma between LLM Capability and OR Problem Complexity
In operations research, most problems can be categorized into a limited number of canonical types, which we treat as the "seed" problems. We deliberately avoid indiscriminate expansion of the StructuredOR dataset because simply modifying parameter values results in highly similar instances. This introduces risks such as test set leakage and model overfitting, reducing the robustness of experimental evaluations.
On the other hand, if we were to diversify the dataset by significantly increasing the size and dimensionality of the problem instances, the resulting questions would contain a large number of numerical values. Current LLMs face fundamental limitations in this regard, as they struggle to memorize and accurately reproduce numerous numerical inputs in a structured and orderly fashion.
For instance, a single parameter such as $\texttt{Cost}[X][A][t]$ with three dimensions—each containing 3 elements—already results in $273^{3}=$ 27 values. As the number of elements in each dimension increases, the parameter space grows exponentially. While this level of complexity is trivial for mathematical solvers, it poses a significant challenge for LLMs, which are not yet capable of accurately reproducing large-scale numerical values in a complete and orderly manner.
@@ A.1 Modeling Data Format Specification
Tables  8910111213141516 define the standardized data format for representing mathematical models. This convention provides a structured and consistent way to label and organize modeling data, ensuring clarity and usability across different tasks and datasets.
@@ A.5 PRM and Preference Model Prompt
A.6 BPP Search Complexity Analysis
The computational complexity of BPP-Search is determined by the number of invocations of the policy model, the pairwise preference model, and the PRM. Specifically, the overall complexity can be decomposed into the following three components:
Here, $h$ denotes the height of the search tree, $b$ represents the beam width, and $n$ is the number of child nodes per parent node.
Policy Model Invocations:
$\mathcal{O}(n\cdot b\cdot(h{-}1)+n)$. This term quantifies the cost incurred by the policy model during the node expansion process across the tree levels.
Pairwise Preference Model Invocations:
$\mathcal{O}(b^{2}\cdot n^{2})$. This cost arises from ranking and selecting candidates in the final layer, where each candidate is compared pairwise with the others.
PRM Invocations:
$\mathcal{O}(n\cdot b\cdot(h{-}2)+n)$. This component reflects the number of calls to the PRM for evaluating the intermediate steps throughout the search.
Since the reward model and the pairwise preference model, typically implemented with 1.5B or 3B parameters, are relatively small compared to the policy model (usually a 70B, 405B parameter model), the primary concern is the time spent calling the policy model.
A.7 PRM, PM Training
Table 17 and Table 18 report the performance of the PRM and the Preference Model on their corresponding test sets after supervised fine-tuning.
Metric Accuracy Precision Recall F1-Score
Value 0.9823 0.9772 0.9868 0.9820
Metric Accuracy Precision Recall F1-Score
Value 0.7560 0.7761 0.7196 0.7468